whatvis the percentage of nitrogen in ammonium sulphate

Grade:12th pass

1 Answers

Arun
6 years ago
(NH4)2SO4 mol mass = (N: 14 x 2) + (H: 1 x 8) + (S: 32 x 1) + (O:16 x 4) = 132g/mol. 

N: = (28 / 132) x 100% = 21.21% Nitrogen. 
H: = (8 / 132) x 100% = 6.06% Hydrogen. 
S: = (32 / 132) x 100% = 24.24% Sulphur. 
O: = (64 / 132) x 100% = 48.49% Oxygen. 

Total: = 100%